OpenStudy (anonymous):

The point P(4,2) lies on the curve y=√x. If Q is the point (x, √x), write an expression that gives the slope of the secant line.

OpenStudy (anonymous):

∆f/∆x= f(x)-f(x)/x-a=> Vx-2/x-4

OpenStudy (anonymous):

something like that.... I guess...

OpenStudy (anonymous):

\[dy/dx=1/2\sqrt{x}\]

OpenStudy (anonymous):

dy/dx is the slope